OpenStudy (anonymous):

Use an Addition or Subtraction Formula to find the exact value of the expression, as demonstrated in Example 1. tan 75°

hartnn (hartnn):

write tan 75 as tan (30+45) and use tan (A+B) = \(\large \frac{tanA+tanB}{1-tanAtanB}\) where A = 30, B=45

hartnn (hartnn):

know what is tan 30 and tan 45 ?

OpenStudy (anonymous):

the tan of 30 is 1/ root 3 tan of 45 is 1

hartnn (hartnn):

that is correct.

OpenStudy (anonymous):

so its root 3 +1/1- root 3 right?

hartnn (hartnn):

\(\large tan (30+45)= \frac{tan30+tan45}{1-tan30tan45}=\frac{1/\sqrt 3+1}{1-1/\sqrt 3}\)

OpenStudy (anonymous):

so 2 - root 3

hartnn (hartnn):

\(\large =\frac{1/\sqrt 3+1}{1-1\sqrt 3}= \frac{1+\sqrt 3 }{\sqrt 3-1}= \frac{(1+\sqrt 3)^2}{3-1}= ?\)

OpenStudy (anonymous):

opps i mean 2 + root 3

hartnn (hartnn):

do u have options for this ?

hartnn (hartnn):

and yes, 2+ root 3 is correct

OpenStudy (anonymous):

no its just an open question from my online hw